OpenStudy (anonymous):

Can someone explain how to find: d\dx tan^-1.

OpenStudy (phi):

Let's say the inverse tan of x gives us an angle theta: \[ \tan^{-1}x = \theta \] take the derivative with respect to "x", of both sides: \[ \frac{d}{dx}\left( \tan^{-1}x\right) = \frac{d \theta}{dx} \] this tells us we want to find d\(\theta\)/dx. How do do that ? if we start with the original problem, and take the tangent of both sides, like so: \[ \tan^{-1}x = \theta \\ x = \tan\theta\] now take the derivative of both sides with respect to theta (you are able to find the derivative of tan, right?) \[ \frac{dx}{d\theta} = \sec^2 \theta = \frac{1}{\cos^2 \theta} \] and if we "flip" both sides: \[ \frac{d\theta}{dx}= \cos^2 \theta \] now we have d\(\theta\)/dx, so we can say \[ \frac{d}{dx}\left( \tan^{-1}x\right) =\frac{d\theta}{dx}= \cos^2 \theta \] that is the answer, except we want the derivative in terms of x and not \(\theta\)

OpenStudy (phi):

to replace the theta, we draw a triangle with angle theta, and sides such that the tangent of theta is x. A reasonable choice of lengths would be:  we see that \[ \cos \theta = \frac{1}{\sqrt{x^2+1}} \\ \] and \[ \cos^2 \theta = \frac{1}{x^2+1} \] and using this result, we get the final answer: \[ \frac{d}{dx} \tan^{-1} x = \frac{1}{x^2+1} \]